# Implicit Regularization of Infinitesimally-perturbed Gradient Descent Toward Low-dimensional Solutions

### v1

@@ 2.1 Efficient Convergence to Approximate SOSPs
We first formalize the intuition behind the PGD algorithm discussed in Section 1.1 by stating its convergence guarantee from [20].
Theorem 1 (Theorem 3 of [20]; informal).
@@
Indeed, if PGD encounters $N_{T}$ approximate strict saddle points along its trajectory, it must apply precisely $N_{T}$ isotropic perturbations, each with radius $\gamma$. Intuitively, a single isotropic perturbation of an approximate strict saddle point $\mathbf{x}$ with radius $\gamma$ can increase its residual norm by $O(\gamma)$. Consequently, the final increase in the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{T}\right\rVert$ caused by perturbations can be as large as $O(\gamma N_{T})$. As shown in [20] and also demonstrated in our analysis, $N_{T}$ can be as large as $O\left(\epsilon^{-3/2}\right)$. Thus, if the perturbation radius is set to $\gamma=\tilde{\Theta}({\epsilon}/L)$, as required by Theorem 1, the final residual norm of PGD scales as $\tilde{O}\left(\epsilon^{-3/2}\cdot{\epsilon}/L\right)=\tilde{O}\left(\epsilon^{-1/2}/L\right)\gg 1$. In other words, ensuring the convergence of PGD to an approximate SOSP may inevitably lead to the divergence of the residual norm in the over-parameterized regime.
To address this challenge, we prove that PGD can efficiently escape approximate strict saddle points using a significantly smaller perturbation radius. The detailed implementation of this variant of PGD, which we call infinitesimally-perturbed gradient descent (IPGD), can be found in Algorithm 1.
@@
2:$F\leftarrow\frac{1}{C}\!\cdot\!\frac{\epsilon^{3/2}}{\sqrt{\rho}\left(\log^{3}\left(\frac{1}{\gamma}\right)+\log^{3}\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)}$, $G\leftarrow\frac{1}{C}\!\cdot\!\frac{\epsilon}{\log^{2}(1/\gamma)}$, $T_{\mathrm{escape}}\leftarrow\frac{C}{\eta\sqrt{\rho\epsilon}}\left(\log\left(\frac{1}{\gamma}\right)+\log\left(\frac{\rho d\Delta_{f}}{\chi\epsilon}\right)\right)$, for a sufficiently large constant $C>0$
3:$T_{\mathrm{noise}}\leftarrow-T_{\mathrm{escape}}-1$
@@
Our first theorem demonstrates that IPGD achieves nearly the same convergence guarantee as PGD (differing only by a logarithmic factor), while requiring an arbitrarily smaller perturbation radius.
@@
Unlike Theorem 1, our result demonstrates that IPGD can efficiently converge to an approximate SOSP using a positive perturbation radius that is significantly smaller than $\tilde{\Theta}(\epsilon/L)$, which is required in Theorem 1. This improvement comes at the cost of increasing the number of iterations only by a factor of $\log^{4}(1/\gamma)$. This implies that the total increase in the residual norm caused by perturbations can be kept small, without significantly impacting the convergence rate.
@@
Analogous to [20], our approach to proving Theorem 2 is based on characterizing the volume of the “stuck region” surrounding an approximate strict saddle point. Specifically, let us assume that $\mathbf{z}$ is a strict saddle point of $f$. [20] defines the stuck region $\mathcal{R}_{\mathrm{stuck}}$ and the escape region $\mathcal{R}_{\mathrm{escape}}$ as follows:
$\mathcal{R}_{\mathrm{escape}}$ :=$:=\{\mathbf{x}_{0}\in\mathbb{B}_{\mathbf{z}}(\gamma):f(\mathbf{x}_{t})-f(\mathbf{x}_{0})<-F_{\mathrm{PGD}},\text{within $t\leq T_{\mathrm{PGD}}$ GD updates}\},$ (3)
$\mathcal{R}_{\mathrm{stuck}}$ :=$:=\{\mathbf{x}_{0}\in\mathbb{B}_{\mathbf{z}}(\gamma):f(\mathbf{x}_{t})-f(\mathbf{x}_{0})\geq-F_{\mathrm{PGD}},\text{within $t\leq T_{\mathrm{PGD}}$ GD updates}\},$ (4)
for appropriately chosen values of $F_{\mathrm{PGD}}$ and $T_{\mathrm{PGD}}$. Evidently, any point in $\mathcal{R}_{\mathrm{escape}}$ successfully escapes the strict saddle point $\mathbf{z}$ within the specified number of GD updates, whereas any point in $\mathcal{R}_{\mathrm{stuck}}$ remains trapped and fails to do so. The key insight from [20] is that $\mathcal{R}_{\mathrm{stuck}}$ forms a narrow band with a “width” proportional to $\epsilon/\sqrt{d}$. Consequently, even if $\mathbf{z}$ lies within this region, a single perturbation with a radius scaling as $\epsilon$ is likely to move the perturbed point into $\mathcal{R}_{\mathrm{escape}}$ (see the right panel of Figure 3). Our main contribution is to demonstrate that the width of $\mathcal{R}_{\mathrm{stuck}}$ can be reduced to $\gamma/\sqrt{d}$, for an arbitrarily small value $\gamma>0$ (potentially much smaller than $\epsilon$), while only requiring $F_{\mathrm{PGD}}$ and $T_{\mathrm{PGD}}$ to scale poly-logarithmically in $1/\gamma$. Specifically, we show that this can be achieved by adjusting $F_{\mathrm{IPGD}}$ and $T_{\mathrm{IPGD}}$ as $F_{\mathrm{IPGD}}\leftarrow F_{\mathrm{PGD}}/\mathrm{poly}\log(1/\gamma)$ and $T_{\mathrm{IPGD}}\leftarrow T_{\mathrm{PGD}}\cdot\mathrm{poly}\log(1/\gamma)$. An immediate implication of this is that a perturbation radius of $\gamma$ is sufficient to guarantee an efficient escape (see the left panel of Figure 3).
@@ 2.2 Implicit Regularization
Indeed, even if the infinitesimal perturbations of IPGD do not cause a significant deviation from the implicit region $\mathcal{M}$, there remains the possibility that the iterates may drift away from this region during the standard GD updates. To characterize this deviation, we provide a fine-grained control over the dynamics of the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$. At a high level, a single GD update can be reformulated as:
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta\nabla f\big{(}\mathbf{x}_{t}\big{)}$ (5)
$=\bm{\mathbf{x}}^{\sharp}_{t}-\eta\nabla f\big{(}\bm{\mathbf{x}}^{\sharp}_{t}\big{)}+\bm{\mathbf{x}}^{\perp}_{t}-\eta\left(\nabla f(\mathbf{x}_{t})-\nabla f\big{(}\bm{\mathbf{x}}^{\sharp}_{t}\big{)}\right)$
@@
The last equality follows from the Hessian-Lipschitzness of $f$. Based on the above equation, the one-step dynamic of the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$ can be controlled by applying the projection $\mathcal{P}^{\perp}_{\mathcal{M}}(\cdot)$ to both sides:
@@
In the above inequality, we used the simple property that, for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, we have $\left\lVert\mathcal{P}^{\perp}_{\mathcal{M}}(\mathbf{x}+\mathbf{y})\right\rVert\leq\left\lVert\mathcal{P}^{\perp}_{\mathcal{M}}(\mathbf{x})\right\rVert+\left\lVert\mathbf{y}\right\rVert$ (see Lemma 16). According to the above decomposition, for sufficiently small residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$, the dynamic of the residual norm is mainly governed by the signal deviation and residual deviation. The favorable landscape properties of $f$ within the implicit domain $\mathcal{M}$ can be leveraged to effectively control the signal deviation. In particular, the signal deviation is completely eliminated if the implicit region $\mathcal{M}$ is closed under GD updates, as formally defined below.
@@
This assumption ensures that the GD iterates remain within the implicit region, provided the initial point lies within that region. While this assumption may appear restrictive at first and is certainly not universally valid for all implicit regions, we demonstrate in the following three examples that it holds for appropriately chosen implicit regions in sparse recovery, low-rank matrix recovery, and, more generally, models involving sparse basis function decomposition.
@@
Consider the sparse recovery problem discussed in Section 1 and its aforementioned implicit region =$\mathcal{M}=\{(\mathbf{u},\mathbf{v}):u_{i}=v_{i}=0\ \text{if}\ \theta^{\star}_{i}=0\}$. For a general loss function $f(\mathbf{u},\mathbf{v})=L(\mathbf{u}\odot\mathbf{v})$, a single GD update can be written as:
@@
Based on the above formula, it is easy to verify that if $(\mathbf{u},\mathbf{v})\in\mathcal{M}$, then $(\mathbf{u}_{+},\mathbf{v}_{+})\in\mathcal{M}$.
Example 2 (Low-rank matrix recovery).
Consider the symmetric low-rank matrix recovery problem discussed in Section 1 and its associated implicit region =$\mathcal{M}=\{\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}:\mathrm{rank}(\mathbf{X})\leq r\}$. For a general loss function $f(\mathbf{X})=L(\mathbf{X}\mathbf{X}^{\top})$, a single GD update can be written as:
$\mathbf{X}_{+}=\mathbf{X}-\eta\nabla f(\mathbf{X})=\left(1-2\nabla L\left(\mathbf{X}\mathbf{X}^{\top}\right)\right)\mathbf{X}.$ (8)
Indeed, it is easy to verify that if $\mathbf{X}\in\mathcal{M}$, then $\mathbf{X}_{+}\in\mathcal{M}$. Moreover, a similar property can be established for other choices of implicit regions (see Section 6), as well as the asymmetric variant of low-rank matrix recovery and its extension to low-rank tensor decomposition.
@@
The recent paper [17] demonstrates that, although the dynamics of GD can vary significantly across different deep neural network models, they tend to exhibit a sparse behavior when projected onto a suitable orthonormal function basis (see [17, Definition 1]). While the explicit characterization of this orthonormal function basis depends on the specific task, its mere existence is sufficient to define the corresponding implicit region as the set of parameters with a sparse representation—characterized by sparse basis coefficients—after projection onto this basis. It can be verified that this implicit region remains closed under GD updates, provided the GD trajectory satisfies a condition known as gradient independence (see [17, Theorem 1]).  [17] illustrates that the gradient independence condition is approximately satisfied across a wide range of deep neural networks.
An arguably more challenging task is to control the residual deviation, which we discuss next. Consider the eigen-decomposition of the Hessian $\nabla^{2}f(\mathbf{x})=\mathbf{V}\mathbf{D}\mathbf{V}^{\top}$, where $\mathbf{V}\in\mathbb{R}^{d\times d}$ is an orthonormal matrix whose columns correspond to the eigenvectors of $\nabla^{2}f(\mathbf{x})$, and $\mathbf{D}\in\mathbb{R}^{d\times d}$ is a diagonal matrix containing the eigenvalues of $\nabla^{2}f(\mathbf{x})$. The decomposition of the Hessian into positive semi-definite (PSD) and negative semi-definite (NSD) components is defined as $\nabla^{2}f(\mathbf{x})=\nabla^{2}_{+}f(\mathbf{x})+\nabla^{2}_{-}f(\mathbf{x})$, where +$\nabla^{2}_{+}f(\mathbf{x})=\mathbf{V}\mathbf{D}_{+}\mathbf{V}^{\top}$ and $\nabla^{2}_{-}f(\mathbf{x})=\mathbf{V}\mathbf{D}_{-}\mathbf{V}^{\top}$. Here $\mathbf{D}_{+}$ is a diagonal matrix whose $(i,i)$-th entry is given by $\max\{0,\mathbf{D}_{ii}\}$. Similarly, $\mathbf{D}_{-}$ is defined as a matrix whose $(i,i)$-th entry is given by $\min\{0,\mathbf{D}_{ii}\}$.
@@
$r_{-}(\mathbf{x})=\begin{cases}\frac{-1}{\left\lVert\bm{\mathbf{x}}^{\perp}\right\rVert^{2}}\left\langle\bm{\mathbf{x}}^{\perp},\nabla^{2}_{+}f(\bm{\mathbf{x}}^{\sharp})\bm{\mathbf{x}}^{\perp}\right\rangle&\text{if $\bm{\mathbf{x}}^{\perp}\not=0;$}\\ 0&\text{if $\bm{\mathbf{x}}^{\perp}=0$}\end{cases},$ (9)
$r_{+}(\mathbf{x})=\begin{cases}\frac{-1}{\left\lVert\bm{\mathbf{x}}^{\perp}\right\rVert^{2}}\left\langle\bm{\mathbf{x}}^{\perp},\nabla^{2}_{-}f(\bm{\mathbf{x}}^{\sharp})\bm{\mathbf{x}}^{\perp}\right\rangle&\text{if $\bm{\mathbf{x}}^{\perp}\not=0$},\\ 0&\text{if $\bm{\mathbf{x}}^{\perp}=0$}.\end{cases}$ (10)
@@
To provide a better intuition on the role of Proposition 4 in our analysis, let us, for simplicity, assume that $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert\leq\tau\leq 2/(T\eta\rho)$ for all $10\leq t\leq T-$ 1.333This assumption is not required in our formal analysis but is introduced here to provide a crisp intuition. Unfolding the dynamics of $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$ for $t=0,1,\dots,T-1$ yields:
@@
Below, we discuss a few key insights from Theorem 5.
@@
The convergence of IPGD to an approximate $\mathcal{M}$-SOSP only requires gradient- and Hessian-Lipschitz properties within an $(\epsilon/L)$-neighborhood of the implicit region $\mathcal{M}$. This is a significant improvement over the global Lipschitz conditions required for PGD and for the convergence of IPGD to a general SOSP (Theorem 2). This relaxation is possible because, by controlling the residual norm via Proposition 4, the trajectory of IPGD can be confined to a small neighborhood around the implicit region.
@@
The convergence of IPGD to an approximate $\mathcal{M}$-SOSP depends on the initial point being close to the implicit region, which corresponds to a small initial residual norm. While it may not always be feasible to determine such an initialization for an arbitrary implicit region $\mathcal{M}$, it is often possible to ensure that $\bm{0}\in\mathcal{M}$. In this case, a small or zero initialization naturally satisfies this condition. Notably, it is straightforward to verify that $\bm{0}\in\mathcal{M}$ holds for the implicit regions discussed in Examples 1, 2 and 3.

### v2

@@ 2.1 Efficient Convergence to Approximate SOSPs
We first formalize the intuition behind the PGD algorithm discussed in Section 1.2 by stating its convergence from [jin2017escape].
Theorem 1 (Theorem 3 of [jin2017escape]; informal).
@@
If PGD encounters $N$ approximate SSPs along its trajectory, it applies $N$ isotropic perturbations, each of radius $\gamma$. A single perturbation around an SSP $\mathbf{x}$ can increase the residual norm by $O(\gamma)$, so the cumulative increase in $\|\bm{\mathbf{x}}^{\perp}_{T}\|$ can reach $O(\gamma N)$. As shown in [jin2017escape] and in our analysis, $N$ can be as large as $O\left(\epsilon^{-3/2}\right)$. Hence, with $\gamma=\tilde{\Theta}(\epsilon/L)$ as required by Theorem 1, the final residual norm scales as $\tilde{O}(\epsilon^{-3/2}\cdot\epsilon/L)=\tilde{O}(\epsilon^{-1/2}/L)\gg 1$. In other words, ensuring the convergence of PGD to an approximate SOSP may inevitably lead to the divergence of the residual norm in the over-parameterized regime.
To address this challenge, we prove that PGD can efficiently escape approximate SSPs using a significantly smaller perturbation radius. The detailed implementation of this variant of PGD, which we call infinitesimally-perturbed gradient descent (IPGD), can be found in Algorithm 1.
@@
[…]
@@
Our first theorem demonstrates that IPGD achieves nearly the same convergence guarantee as PGD (differing only by a logarithmic factor), while requiring a smaller perturbation radius.
@@
Unlike Theorem 1, our result demonstrates that IPGD can efficiently converge to an approximate SOSP using a perturbation radius that is significantly smaller than $\tilde{\Theta}(\epsilon/L)$, which is required in Theorem 1. This improvement comes at the cost of increasing the number of iterations only by a factor of $\log^{4}(1/\gamma)$. This implies that the total increase in the residual norm caused by perturbations can be kept small, without significantly impacting the convergence rate.
@@
Analogous to [jin2017escape], our approach to proving Theorem 2 is based on characterizing the volume of the “stuck region” surrounding an approximate SSP. Specifically, let us assume that $\mathbf{z}$ is a SSP of $f$. [jin2017escape] defines the stuck region $\mathcal{R}_{\mathrm{stuck}}$ and the escape region $\mathcal{R}_{\mathrm{escape}}$ as follows:
$\mathcal{R}_{\mathrm{escape}}$ $:=\{\mathbf{x}_{0}\in\mathbb{B}_{\mathbf{z}}(\gamma):f(\mathbf{x}_{t})-f(\mathbf{x}_{0})<-F_{\mathrm{PGD}},\text{within $t\leq T_{\mathrm{PGD}}$ GD updates}\},$ (5)
$\mathcal{R}_{\mathrm{stuck}}$ $:=\{\mathbf{x}_{0}\in\mathbb{B}_{\mathbf{z}}(\gamma):f(\mathbf{x}_{t})-f(\mathbf{x}_{0})\geq-F_{\mathrm{PGD}},\text{within $t\leq T_{\mathrm{PGD}}$ GD updates}\},$ (6)
for appropriately chosen values of $F_{\mathrm{PGD}}$ and $T_{\mathrm{PGD}}$. Evidently, any point in $\mathcal{R}_{\mathrm{escape}}$ successfully escapes the strict saddle point $\mathbf{z}$ within the specified number of GD updates, whereas any point in $\mathcal{R}_{\mathrm{stuck}}$ remains trapped and fails to do so. The key insight from [jin2017escape] is that $\mathcal{R}_{\mathrm{stuck}}$ forms a narrow band with a “width” proportional to $\epsilon/\sqrt{d}$. Consequently, even if $\mathbf{z}$ lies within this region, a single perturbation with a radius scaling as $\epsilon$ is likely to move the perturbed point into $\mathcal{R}_{\mathrm{escape}}$ (see the right panel of Figure 3). Our main contribution is to demonstrate that the width of $\mathcal{R}_{\mathrm{stuck}}$ can be reduced to $\gamma/\sqrt{d}$, for an arbitrarily small value $\gamma>0$ (potentially much smaller than $\epsilon$), while only requiring $F_{\mathrm{PGD}}$ and $T_{\mathrm{PGD}}$ to scale poly-logarithmically in $1/\gamma$. Specifically, we show that this can be achieved by adjusting $F_{\mathrm{IPGD}}$ and $T_{\mathrm{IPGD}}$ as $F_{\mathrm{IPGD}}\leftarrow F_{\mathrm{PGD}}/\mathrm{poly}\log(1/\gamma)$ and $T_{\mathrm{IPGD}}\leftarrow T_{\mathrm{PGD}}\cdot\mathrm{poly}\log(1/\gamma)$. An immediate implication of this is that a perturbation radius of $\gamma$ is sufficient to guarantee an efficient escape (see the left panel of Figure 3).
@@ 2.2 Implicit Regularization
Even if the infinitesimal perturbations of IPGD do not cause a significant deviation from the implicit region $\mathcal{M}$, there remains the possibility that the iterates may drift away from this region during the standard GD updates. To characterize this deviation, we provide a control over the dynamics of the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$. At a high level, a single GD update can be reformulated as:
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta\nabla f\big(\mathbf{x}_{t}\big)=\bm{\mathbf{x}}^{\sharp}_{t}-\eta\nabla f\big(\bm{\mathbf{x}}^{\sharp}_{t}\big)+\bm{\mathbf{x}}^{\perp}_{t}-\eta\left(\nabla f(\mathbf{x}_{t})-\nabla f\big(\bm{\mathbf{x}}^{\sharp}_{t}\big)\right)$ (7)
@@
The last equality follows from the Hessian-Lipschitzness of $f$. Therefore, the one-step dynamic of the residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$ can be controlled by applying the projection $\mathcal{P}^{\perp}_{\mathcal{M}}(\cdot)$ to both sides:
@@
In the above inequality, we used the simple property that, for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, we have $\left\lVert\mathcal{P}^{\perp}_{\mathcal{M}}(\mathbf{x}+\mathbf{y})\right\rVert\leq\left\lVert\mathcal{P}^{\perp}_{\mathcal{M}}(\mathbf{x})\right\rVert+\left\lVert\mathbf{y}\right\rVert$ (see Lemma 6). According to the above decomposition, for sufficiently small residual norm $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$, the dynamic of the residual norm is mainly governed by the signal deviation and residual deviation. The favorable landscape properties of $f$ within the implicit region $\mathcal{M}$ can be leveraged to effectively control the signal deviation. In particular, the signal deviation is completely eliminated if the implicit region $\mathcal{M}$ is closed under GD updates.
@@
This assumption ensures that the GD iterates remain within the implicit region, provided the initial point lies within that region. While this assumption may appear restrictive at first and is certainly not universally valid for all choices of $\mathcal{M}=\varphi^{-1}(\mathcal{D})$, we demonstrate in the following three examples that it holds, possibly with minor modifications, in low-rank matrix recovery, sparse recovery, and, more generally, models involving sparse basis function decomposition.
Example 1 (Low-rank matrix recovery).
Consider the symmetric low-rank matrix recovery problem discussed in Section 1 with the feasible region $\mathcal{D}=\{\mathbf{\Theta}:\mathbf{\Theta}\succeq 0,\mathrm{rank}(\mathbf{\Theta})\leq r\}$. With reparameterization map $\varphi:\mathbb{R}^{n\times r^{\prime}}\to\mathbb{R}^{n\times n}$ with $r^{\prime}\geq r$ and $\varphi(\mathbf{X})=\mathbf{X}\mathbf{X}^{\top}$, the associated implicit region is defined as $\mathcal{M}=\varphi^{-1}(\mathcal{D})=\{\mathbf{X}\in\mathbb{R}^{n\times r^{\prime}}:\mathrm{rank}(\mathbf{X})\leq r\}$. For a general function $f(\mathbf{X})=L(\mathbf{X}\mathbf{X}^{\top})$, a single GD update can be written as:
$\mathbf{X}_{+}=\mathbf{X}-\eta\nabla f(\mathbf{X})=\left(\mathbf{I}-2\eta\nabla L\left(\mathbf{X}\mathbf{X}^{\top}\right)\right)\mathbf{X}.$ (9)
It is easy to verify that if $\mathbf{X}\in\mathcal{M}$, then $\mathbf{X}_{+}\in\mathcal{M}$. A similar property can be established for the asymmetric low-rank matrix recovery and its extension to low-rank tensor decomposition.
In many applications, however, it may be preferable to work with a smaller implicit region $\mathcal{M}\subset\varphi^{-1}(\mathcal{D})$, as requiring 1 to hold over the entire set $\varphi^{-1}(\mathcal{D})$ could be too restrictive. Importantly, the reparameterization map $\varphi$ is typically not bijective. By carefully choosing a smaller implicit region $\mathcal{M}$, we can guarantee that 1 holds, while still ensuring that $\varphi(\mathcal{M})=\mathcal{D}$. The following example illustrates this idea in the context of sparse recovery.
@@
Consider the sparse recovery discussed in Section 1 with the feasible region $\mathcal{D}=\{\bm{\theta}\in\mathbb{R}^{n}:\bm{\theta}_{i}=0,\ \text{if $\bm{\theta}^{\star}_{i}=0$}\}$. Under the reparameterization map $\varphi(\mathbf{u},\mathbf{v})=\mathbf{u}\odot\mathbf{v}$, the implicit region is $\mathcal{M}=\varphi^{-1}(\mathcal{D})=\{(\mathbf{u},\mathbf{v}):\mathbf{u}_{i}=0\ \text{or}\ \mathbf{v}_{i}=0\ \text{if}\ \bm{\theta}^{\star}_{i}=0\}$. For a general function $f(\mathbf{u},\mathbf{v})=L(\mathbf{u}\odot\mathbf{v})$, a single GD update takes the form
@@
For this problem, $(\mathbf{u},\mathbf{v})\in\mathcal{M}$ alone does not guarantee that $(\mathbf{u}_{+},\mathbf{v}_{+})\in\mathcal{M}$. This issue can be resolved by restricting $\mathcal{M}$ to the set of balanced vectors, defined as $\mathcal{M}_{\text{balanced}}=\{(\mathbf{u},\mathbf{v}):\mathbf{u}_{i}=\mathbf{v}_{i}=0\ \text{if}\ \bm{\theta}^{\star}_{i}=0\}$. This restricted implicit region continues to satisfy $\varphi(\mathcal{M}_{\text{balanced}})=\mathcal{D}$, while also ensuring that $(\mathbf{u},\mathbf{v})\in\mathcal{M}_{\text{balanced}}$ implies $(\mathbf{u}_{+},\mathbf{v}_{+})\in\mathcal{M}_{\text{balanced}}$. In Section 7, we empirically show that IPGD exhibits implicit regularization toward solutions in $\mathcal{M}_{\text{balanced}}$.
@@
The recent paper [ma2023behind] demonstrates that, although the dynamics of GD can vary significantly across different deep neural network models, they tend to exhibit a sparse behavior when projected onto a suitable orthonormal function basis (see [ma2023behind, Definition 1]). While the explicit characterization of this orthonormal function basis depends on the specific task, its mere existence is sufficient to define the corresponding implicit region as the set of parameters with a sparse representation—characterized by sparse basis coefficients—after projection onto this basis. It can be verified that this implicit region remains closed under GD updates, provided the GD trajectory satisfies a condition known as gradient independence (see [ma2023behind, Theorem 1]).  [ma2023behind] illustrates that the gradient independence condition is approximately satisfied across a wide range of deep neural networks.
An arguably more challenging task is to control the residual deviation, which we discuss next. Consider the eigen-decomposition of the Hessian $\nabla^{2}f(\mathbf{x})=\mathbf{V}\mathbf{D}\mathbf{V}^{\top}$, where $\mathbf{V}\in\mathbb{R}^{d\times d}$ is an orthonormal matrix whose columns correspond to the eigenvectors of $\nabla^{2}f(\mathbf{x})$, and $\mathbf{D}\in\mathbb{R}^{d\times d}$ is a diagonal matrix containing the eigenvalues of $\nabla^{2}f(\mathbf{x})$. The decomposition of the Hessian into positive semi-definite (PSD) and negative semi-definite (NSD) components is defined as $\nabla^{2}f(\mathbf{x})=\nabla^{2}_{+}f(\mathbf{x})+\nabla^{2}_{-}f(\mathbf{x})$, where $\nabla^{2}_{+}f(\mathbf{x})=\mathbf{V}\mathbf{D}_{+}\mathbf{V}^{\top}$ and $\nabla^{2}_{-}f(\mathbf{x})=\mathbf{V}\mathbf{D}_{-}\mathbf{V}^{\top}$. Here $\mathbf{D}_{+}$ is a diagonal matrix whose $(i,i)$-th entry is given by $\max\{0,\mathbf{D}_{ii}\}$. Similarly, $\mathbf{D}_{-}$ is defined as a matrix whose $(i,i)$-th entry is given by $\min\{0,\mathbf{D}_{ii}\}$.
@@
$r_{-}(\mathbf{x})\!=\!\begin{cases}\frac{-1}{\left\lVert\bm{\mathbf{x}}^{\perp}\right\rVert^{2}}\left\langle\bm{\mathbf{x}}^{\perp},\nabla^{2}_{+}f(\bm{\mathbf{x}}^{\sharp})\bm{\mathbf{x}}^{\perp}\right\rangle&\text{if $\bm{\mathbf{x}}^{\perp}\not=0;$}\\ 0&\text{if $\bm{\mathbf{x}}^{\perp}=0$}\end{cases},r_{+}(\mathbf{x})\!=\!\begin{cases}\frac{-1}{\left\lVert\bm{\mathbf{x}}^{\perp}\right\rVert^{2}}\left\langle\bm{\mathbf{x}}^{\perp},\nabla^{2}_{-}f(\bm{\mathbf{x}}^{\sharp})\bm{\mathbf{x}}^{\perp}\right\rangle&\text{if $\bm{\mathbf{x}}^{\perp}\not=0$},\\ 0&\text{if $\bm{\mathbf{x}}^{\perp}=0$}.\end{cases}$ (11)
@@
To provide a better intuition on the role of Proposition 1 in our analysis, let us, for simplicity, assume that $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert\leq\tau\leq 2/(T\eta\rho)$ for all $0\leq t\leq T-1$. We note that this assumption is not required in our formal analysis but is introduced here to provide a crisp intuition. Unfolding the dynamics of $\left\lVert\bm{\mathbf{x}}^{\perp}_{t}\right\rVert$ for $t=0,1,\dots,T-1$ using (13) yields:
@@
Below, we discuss a few key insights.
@@
The convergence of IPGD to an approximate $\mathcal{M}$-SOSP only requires gradient- and Hessian-Lipschitz properties within an $(\epsilon/L)$-neighborhood of the implicit region $\mathcal{M}$. This is a significant improvement over the global Lipschitz conditions required for PGD and for the convergence of IPGD to a general SOSP (Theorem 2), especially when $\mathcal{M}$ is low-dimensional. This relaxation is possible because, by controlling the residual norm via Proposition 1, the trajectory of IPGD can be confined to a small neighborhood around the implicit region.
@@
The convergence of IPGD to an approximate $\mathcal{M}$-SOSP depends on the initial point being close to the implicit region, which corresponds to a small initial residual norm. As we mentioned earlier, while it may not always be feasible to determine such an initialization for an arbitrary implicit region $\mathcal{M}$, it is often possible to ensure that $\bm{0}\in\mathcal{M}$. In this case, a small or zero initialization naturally satisfies this condition. Notably, it is straightforward to verify that $\bm{0}\in\mathcal{M}$ holds for the implicit regions discussed in Examples 2, 1 and 3.